# Reliability, and Global Privacy

### v3

@@ I-B Calculation of $\bar{\nu}_{f^{\prime}}$ and $\mathbb{E}[(\nu_{f^{\prime}}^{t})^{2}]$
Thus, $\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})=\nu_{f^{\prime}}^{t}\bm{I}_{D}$, where
$\nu_{f^{\prime}}^{t}$ $=\sum_{l=1}^{L}\left(\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\alpha_{m,l}\right)^{2}$ (128a)
@@
$\bar{\nu}_{f^{\prime}}=\mathbb{E}\left[\nu_{f^{\prime}}^{t}\right]=\sum_{l=1}^{L}\bm{\alpha}_{l}^{\top}\left(\sum_{k_{1},k_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}\tau_{k_{1}}^{t}\tau_{k_{2}}^{t}\Sigma_{k}\right)\bm{\alpha}_{l},$
@@
Next, let us calculate $\mathbb{E}\left[(\nu_{f^{\prime}}^{t})^{2}\right]$. The (128a) can be rewritten as
$\nu_{f^{\prime}}^{t}=\sum_{l=1}^{L}\left(\sum_{k_{1},m_{1}}c_{f^{\prime},k_{1}}g_{k_{1},m_{1}}\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}\alpha_{m_{1},l}\right)^{2}$ (131a)
@@
$\nu_{f^{\prime}}^{t}=\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}H_{m_{1},m_{2}}t(k_{1},m_{1})t(k_{2},m_{2}).$ (132)
@@
$(\nu_{f^{\prime}}^{t})^{2}=\hskip-5.69054pt\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}\sum_{k_{3},m_{3}}\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
@@
$\mathbb{E}\left[(\nu_{f^{\prime}}^{t})^{2}\right]\hskip-2.84526pt=\hskip-5.69054pt\sum_{k_{1},m_{1}}\hskip-1.42262pt\sum_{k_{2},m_{2}}\hskip-1.42262pt\sum_{k_{3},m_{3}}\hskip-1.42262pt\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$

### v4

@@ G-B Calculation of $\bar{\nu}_{F_{f^{\prime}}}$ and $\mathbb{E}[(\nu_{F_{f^{\prime}}}^{t})^{2}]$
Thus, $\mathrm{var}(\bm{N}_{F_{f^{\prime}}}^{t})=\nu_{F_{f^{\prime}}}^{t}\bm{I}_{D}$, where
$\nu_{F_{f^{\prime}}}^{t}$ $=\sum_{l=1}^{L}\left(\sum_{k=1}^{K}c_{f^{\prime},k}\tau_{k}^{t}\sum_{m=1}^{K}\tau_{k,m}^{t}g_{k,m}\alpha_{m,l}\right)^{2}$ (106a)
@@
$\bar{\nu}_{F_{f^{\prime}}}=\mathbb{E}\left[\nu_{F_{f^{\prime}}}^{t}\right]=\sum_{l=1}^{L}\bm{\alpha}_{l}^{\top}\left(\sum_{k_{1},k_{2}}c_{f^{\prime},k_{1}}c_{f^{\prime},k_{2}}\tau_{k_{1}}^{t}\tau_{k_{2}}^{t}\Sigma_{k}\right)\bm{\alpha}_{l},$ (107)
@@
Next, let us calculate $\mathbb{E}\left[(\nu_{F_{f^{\prime}}}^{t})^{2}\right]$. The (106a) can be rewritten as
$\nu_{F_{f^{\prime}}}^{t}=\sum_{l=1}^{L}\left(\sum_{k_{1},m_{1}}c_{f^{\prime},k_{1}}g_{k_{1},m_{1}}\tau_{k_{1}}^{t}\tau_{k_{1},m_{1}}^{t}\alpha_{m_{1},l}\right)^{2}$ (109a)
@@
$\nu_{F_{f^{\prime}}}^{t}=\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}H_{m_{1},m_{2}}t(k_{1},m_{1})t(k_{2},m_{2}).$ (110)
@@
$(\nu_{F_{f^{\prime}}}^{t})^{2}=\hskip-5.69054pt\sum_{k_{1},m_{1}}\sum_{k_{2},m_{2}}\sum_{k_{3},m_{3}}\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
@@
$\mathbb{E}\left[(\nu_{F_{f^{\prime}}}^{t})^{2}\right]\hskip-2.84526pt=\hskip-5.69054pt\sum_{k_{1},m_{1}}\hskip-1.42262pt\sum_{k_{2},m_{2}}\hskip-1.42262pt\sum_{k_{3},m_{3}}\hskip-1.42262pt\sum_{k_{4},m_{4}}A_{k_{1},m_{1}}^{f^{\prime}}A_{k_{2},m_{2}}^{f^{\prime}}A_{k_{3},m_{3}}^{f^{\prime}}A_{k_{4},m_{4}}^{f^{\prime}}$
@@
Appendix H Complete Proof of Theorem VI.1
Under Assumption VI.1, it holds that
Next, we bound the two terms in (114), respectively.
H-A Bounding the First Term
For the first term, we have
$=\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$
$=-\eta\hskip-2.84526pt\sum_{i_{s}=1}^{R_{t_{s}}I}\hskip-2.84526pta_{i_{s}}\mathbb{E}\hskip-2.84526pt\left[\hskip-2.84526pt\left\langle\hskip-2.84526pt\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1}),\frac{1}{K}\hskip-2.84526pt\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})\hskip-2.84526pt\right\rangle\hskip-2.84526pt\right]$ (115a)
$\leq-\frac{1}{2}\eta\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\hskip-2.84526pt\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\hskip-2.84526pt\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}_{k}^{t_{s}-1})\hskip-2.84526pt-\hskip-2.84526pt\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt-\frac{1}{2}\eta\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\frac{1}{K}\sum_{k=1}^{K}\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s},i_{s}-1})\right\lVert^{2}\right]$ (115b)
$\leq-\frac{1}{2}\eta\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\left[\left\lVert\nabla\mathcal{L}(\bm{\Theta}^{t_{s}-1})\right\lVert^{2}\right]$
$\hskip 14.22636pt+\frac{1}{2}\eta\frac{G^{2}}{K}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\bm{\Theta}_{k}^{t_{s}-1}-\bm{\Theta}^{t_{s},i_{s}-1}\right\lVert^{2}\right],$ (115c)
where (115a) follows Assumption VI.2, (115b) is due to the convexity of $\ell_{2}$-norm, and (115c) is due to Assumption VI.1.
H-B Bounding the Second Term
For the second term, it holds that
$=\frac{G}{2}\mathbb{E}\left[\left\lVert-\eta\frac{1}{K}\sum_{k=1}^{K}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\right\rVert^{2}\right]$
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}\mathbb{E}\left[\left\lVert\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\right\rVert^{2}\right]$ (116a)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\hskip-1.42262pt\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})\right\rVert^{2}\right]$ (116b)
$=\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\Bigg(\mathbb{E}\bigg[\bigg\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1},\bm{\xi}_{k}^{t_{s},i_{s}})$
$\hskip 14.22636pt-\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})\bigg\rVert^{2}\bigg]+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})\right\rVert^{2}\right]\Bigg)$ (116c)
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\Bigg(\sigma^{2}+\mathbb{E}\bigg[\bigg\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}_{k}^{t_{s},i_{s}-1})$
$\leq\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}^{2}\lVert\bm{a}\rVert_{1}^{2}\sigma^{2}+\frac{\eta^{2}G}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\cdot$
$\bigg(G^{2}\mathbb{E}\bigg[\bigg\lVert\bm{\Theta}_{k}^{t_{s},i_{s}-1}-\bm{\Theta}^{t_{s}-1}\bigg\rVert^{2}\bigg]+\mathbb{E}\left[\left\lVert\nabla\mathcal{L}_{k}(\bm{\Theta}^{t_{s}-1})\right\rVert^{2}\right]\Bigg)$ (116e)
$+\frac{\eta^{2}G^{3}}{2}\frac{1}{K}\sum_{k=1}^{K}R_{t_{s}}\lVert\bm{a}\rVert_{1}\sum_{i_{s}=1}^{R_{t_{s}}I}a_{i_{s}}\mathbb{E}\bigg[\bigg\lVert\bm{\Theta}_{k}^{t_{s},i_{s}-1}-\bm{\Theta}^{t_{s}-1}\bigg\rVert^{2}\bigg],$ (116f)
where (116a) and (116b) use Jensen’s Inequality, (116c) is due to unbiased estimation, (116d) follows Assumption VI.2, (116e) follows Assumption VI.1, and (116f) follows Assumption VI.3.
H-C Immediate Result
Substitute (115c) and (116f) into (114), then substitute Lemma VI.2, we obtain the immediate result for one round,
Using Assumption VI.3, this further becomes
H-D Average over Successful Rounds
H-E Relaxation with Probability Guarantee
By (127), (120) can be further bounded as follows with probability $0.9974$,
Substitute $e_{R},\nu_{R},e_{R^{2}},\nu_{R^{2}}$, we get
$\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]\hskip-1.42262pt\leq\hskip-1.42262pt\frac{1}{T_{s}}\hskip-2.84526pt\sum_{t_{s}=1}^{T_{s}}\hskip-2.84526ptR_{t_{s}}\mathbb{E}\left[\lVert\Delta\mathcal{L}(\bm{\Theta}^{t_{s}-1})\rVert^{2}\right]$
Appendix I Renewal Theory
According to the algorithm’s stopping rule, training terminates at the first round in which the global model is successful and the cumulative number of training rounds reaches the threshold $T$. Hence, the total number of the actually executed training rounds is
$T_{s}=\min\left\{n:\sum_{t_{s}=1}^{n}R_{t_{s}}\geq T\right\},$ (123)
$e_{R^{2}}=\mathbb{E}[R_{t_{s}}^{2}]=\hskip-5.69054pt\sum_{R_{t_{s}}=1}^{+\infty}\hskip-2.84526ptR_{t_{s}}^{2}P_{O}^{R_{t_{s}}-1}(1-P_{O})=\hskip-2.84526pt\frac{1-P_{O}}{P_{O}}\mathrm{Li}_{-2}(P_{O}),$ (124a)
$\nu_{R^{2}}=\mathrm{var}(R_{t_{s}}^{2})=\mathbb{E}[R_{t_{s}}^{4}]-e_{R^{2}}^{2}$
According to the strong law of large numbers, the empirical average of any function $f(R)$ over the first $T_{s}$ renewals converges to its expected value almost surely, i.e., $\frac{1}{T_{s}}\sum_{t_{s}=1}^{T_{s}}R_{t_{s}}^{2}\rightarrow\mathbb{E}[R_{t_{s}}^{2}]$. According to CLT,